# Few-Shot Learning from Gigapixel Images via Hierarchical Vision-Language Alignment and Modeling

### v4

@@ 4.1 Experimental Settings
ABMIL [27] 84.30 ±2.22 90.97 ±0.60 84.28 ±2.21 81.04 ±3.05 87.50 ±5.38 80.93 ±3.04 88.43 ±1.95 96.17 ±0.76 86.95 ±2.33
DSMIL [33] 85.83 ±2.78 94.23 ±1.20 85.76 ±2.84 82.08 ±3.92 89.91 ±5.46 81.99 ±3.89 91.95 ±1.95 98.20 ±0.23 90.87 ±2.00
CLAM-SB [38] 85.83 ±4.25 93.19 ±2.39 85.80 ±4.29 82.29 ±7.42 90.70 ±6.73 82.24 ±7.41 92.11 ±0.52 98.17 ±0.33 90.76 ±0.85
CLAM-MB [38] 86.92 ±3.39 94.01 ±2.16 86.91 ±3.40 81.88 ±4.82 90.41 ±5.14 81.84 ±4.81 91.42 ±1.13 98.15 ±0.22 89.96 ±1.11
TransMIL [47] 85.90 ±3.36 93.38 ±2.11 85.88 ±3.36 82.50 ±5.37 89.69 ±4.54 82.38 ±5.36 89.27 ±2.34 97.75 ±0.69 87.66 ±2.95
DTFD-MIL [54] 88.40 ±3.54 95.36 ±1.52 88.37 ±3.56 83.54 ±3.86 91.22 ±3.39 83.48 ±3.83 91.65 ±1.44 97.99 ±0.09 90.38 ±1.52
WiKG [34] 82.24 ±3.13 91.17 ±1.62 82.15 ±3.21 79.58 ±6.17 87.42 ±6.54 79.44 ±6.39 89.73 ±2.37 97.65 ±0.67 87.84 ±3.12
ViLa-MIL [48] 83.08 ±3.63 91.10 ±2.43 83.04 ±3.64 77.08 ±6.69 87.03 ±8.01 76.98 ±6.73 89.27 ±2.32 97.48 ±0.79 87.91 ±2.88
MSCPT [22] 80.06 ±5.20 88.06 ±6.28 79.95 ±5.24 79.79 ±8.22 87.33 ±6.78 79.69 ±8.21 92.03 ±1.52 98.03 ±0.35 90.89 ±1.94
FOCUS [22] 85.32 ±2.54 93.43 ±1.45 85.24 ±2.60 82.50 ±5.57 90.10 ±4.50 82.20 ±5.77 91.57 ±1.14 98.13 ±0.54 90.21 ±1.37
@@ 4.2 Main Results
We adopt the 16-shot evaluation setting as our main experimental setup, following [48, 22, 41], and report results across three TCGA datasets (NSCLC, BRCA, RCC) and three vision-language pathology foundation models (PLIP, QuiltNet, CONCH). As shown in Table 1, HiVE-MIL consistently achieves the best performance across all settings, outperforming both traditional and VLM-based MIL methods, including their single-scale and multi-scale variants. Compared to the state-of-the-art baselines, HiVE-MIL achieves significant improvements across all datasets, with gains of up to +4.37% in ACC, +4.22% in AUC, and +4.14% in macro F1 on BRCA, and notable margins on NSCLC (up to +3.27% ACC, +3.28% AUC, +3.26% F1) and RCC (up to +1.07% ACC, +0.85% AUC, +1.64% F1). Even when paired with CONCH, the best pathology VLM to date, HiVE-MIL maintains consistent improvements across all datasets. Traditional MIL methods are based solely on visual features and large WSI labels, making them ineffective in the FSWC setting. Existing VLM-MIL methods, including multi-scale approaches, fail to capture hierarchical interactions across scales and do not effectively align modalities within the same scale. In contrast, HiVE-MIL explicitly models cross-scale hierarchies and intra-scale modality alignments, leading to superior performance.
@@ 5 Discussion
Acknowledgement.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) under grant numbers RS-2022-NR068758 and NRF-2025-25432820.
@@ A.1 Dataset Descriptions
Following [48], we split each dataset into training, validation, and test sets using a 4:3:3 ratio. For a few-shot evaluation, we repeat the random-splitting process five times and report averaged results. Due to the class imbalance in the TCGA BRCA (approximately 20:80), we balance the test set during sampling to avoid misleading final metric results. All details regarding the data splits used in the 4-shot, 8-shot, and 16-shot settings are provided in the splits directory of our GitHub repository (linked in the Abstract).
@@ A.2 Hierarchical Patch Extraction
Existing MIL methods [33, 48, 22, 41] typically process patch features at different scales independently, without preserving any hierarchical structure between them. This design hinders the ability to model structured cross-scale interactions, which are essential in tasks such as WSI classification, where coarse tissue-level patterns and fine-grained cellular features are often interdependent. To address this, we propose a hierarchical extraction pipeline with an explicit 1-to-16 mapping across scales: each low-scale ($5\times$) patch corresponds to a spatial region at the high scale ($20\times$), which is uniformly subdivided into a fixed $4\times 4$ grid, yielding 16 high-scale patches. This structured mapping ensures that each high-scale patch can be directly and uniquely traced back to its low-scale parent, enabling consistent alignment and hierarchical reasoning across scales.
@@ B.1 Class-Discriminative Relevance
We evaluate the faithfulness of the generated descriptions using the LLM-as-a-Judge [16], where LLMs are prompted to infer class labels based solely on the generated descriptions. We employ two evaluators: GPT-4o and GPT o3, to assess descriptions generated by several LLMs: Deep Seek R1 [18], Grok 3 [53], Gemini 2.5 Pro [3], and GPT-4o [1]. Although not perfect, the number of correct predictions is still consistently high, suggesting that the generated text is generally reliable while capturing class-discriminative signals.
@@
DeepSeek R1 [18] 32 (8/24) 28 (8/20) 28 (8/20) 31 (8/23) 28 (8/20)
Grok 3 [53] 32 (8/24) 28 (7/21) 28 (8/20) 32 (8/24) 21 (8/23)
Gemini 2.5 Pro [53] 32 (8/24) 30 (8/22) 29 (7/22) 30 (7/23) 32 (8/24)
GPT-4o [1] 32 (8/24) 30 (8/22) 30 (7/23) 31 (8/23) 32 (8/24)
@@ B.2 Morphology- and Concept-Level Semantic Validation
Table B.2 reports intra- and inter-category text similarity scores, where intra denotes similarities among descriptions of the same morphology or concept, and inter denotes similarities across different ones. Similarity is computed separately at each scale and jointly across both scales using the CONCH text encoder [39] and cosine similarity. These results demonstrate that the descriptions encode highly discriminative and consistent semantics across both morphological groups and scales, and across different LLMs.
@@
DeepSeek R1 [18] 1.000 (0.157) 0.515 (0.164) 0.456 (0.158) 1.000 (0.165) 0.438 (0.141) 0.412 (0.138)
Grok 3 [53] 1.000 (0.123) 0.511 (0.170) 0.434 (0.160) 1.000 (0.186) 0.490 (0.214) 0.415 (0.212)
Gemini 2.5 Pro [3] 1.000 (0.111) 0.558 (0.165) 0.478 (0.150) 1.000 (0.206) 0.491 (0.127) 0.426 (0.142)
GPT-4o [1] 1.000 (0.066) 0.516 (0.114) 0.454 (0.109) 1.000 (0.163) 0.477 (0.108) 0.413 (0.125)
@@ Appendix E Message Passing Details
The message passing process is as follows. During neighbor aggregation, information is collected from neighboring nodes connected via a specific relation type (i.e., edge type). These include intra-scale relations, such as valid patch-text connections at low or high scale, and hierarchical relations, such as connections between low-scale and high-scale patches, or between low-scale and high-scale text. The aggregated neighbor features are then concatenated with the node’s own features and passed through a relation-specific linear transformation (i.e., weight matrix). For intra-scale relations, we adopt the GraphSAGE [21] operator via the HeteroConv module, whereas for hierarchical relations, we employ a Modality-Scale Attention (MSA) mechanism that accounts for both scale directionality and modality type. Finally, a non-linear activation function (ReLU) is applied to enhance representational capacity. By decoupling both aggregation and transformation parameters by relation type, our model effectively encodes heterogeneous and hierarchical information.
@@ Appendix H Implementation Details
We set the logit scaling factor $\gamma$ to 4.5871 (PLIP [25]), 4.6052 (QuiltNet [26]), and 4.0315 (CONCH [39]), using the values provided by each pre-trained VLM. We retain the top-2 patches at $5\times$ and the top-100 patches at $20\times$ scales to compute the final logits. Pooling- and MIL-based methods use only the image encoder, while VLM-based MIL methods leverage both image and text encoders. Most methods operate on single-scale $20\times$ patches, whereas DSMIL [33], ViLa-MIL [48], MSCPT [22], and HiVE-MIL (Ours) utilize multi-scale inputs from both $5\times$ and $20\times$ patches.
Our implementation is based on the official ViLa-MIL codebase [48] and all baselines as well as HiVE-MIL are implemented within this unified framework to ensure fair comparison. Our graph-based modules are implemented using PyTorch Geometric [13]. All experiments are conducted on Ubuntu 20.04.6 using a workstation equipped with two NVIDIA A100 GPUs (40 GB each); however, only one GPU is used for training each model. Complete package versions and dependencies are listed in the requirements.txt file available in our GitHub repository (linked in the Abstract).
@@ Appendix I Baselines
  • PLIP [25]: A vision-language pathology model pretrained on OpenPath, a large dataset of 208,414 image–text pairs curated from public platforms like medical Twitter, enabling strong zero-shot classification and case retrieval via image or language queries.
  • QuiltNet [26]: A vision-language model pretrained on Quilt, a 1M-pair dataset curated from YouTube, medical Twitter, and academic sources, enabling strong zero-shot classification and cross-modal retrieval across diverse histopathology datasets.
  • CONCH [39]: A recent state-of-the-art vision-language pathology foundation model, pretrained on 1.17 million image–caption pairs from diverse biomedical sources, enabling strong zero-shot and transferable performance across 14 benchmarks spanning classification, segmentation, captioning, and retrieval tasks.
@@
  • ABMIL [27]: An attention-based pooling framework that assigns adaptive weights to instances for more informative bag-level aggregation.
  • DSMIL [33]: A dual-stream MIL model where one stream identifies a critical instance via max pooling, while the other aggregates instances by distance-weighted similarity.
  • CLAM [38]: A weakly-supervised MIL framework that combines attention pooling with instance-level clustering to enhance interpretability and slide-level prediction.
  • TransMIL [47]: A transformer-based MIL model that captures inter-instance correlations by modeling both spatial and morphological relationships among patches.
  • DTFD-MIL [54]: A double-tier MIL framework that introduces pseudo-bags to improve training diversity and derives instance probabilities under an attention-based setting.
  • WiKG [34]: A graph-based MIL approach that constructs WSIs as knowledge graphs with directed edges and updates features using knowledge-aware attention.
  • ViLa-MIL [48]: A dual-scale VLM-MIL model generates LLM-based prompts and uses a prototype-guided image decoder and a context-guided text decoder.
  • MSCPT [22]: A multi-scale VLM MIL model that integrates multi-scale visual inputs with LLM-generated prompts through graph-based reasoning and cross-scale aggregation.
  • FOCUS [19]: A single-scale VLM-MIL model that integrates pathology foundation models with language priors for focused analysis of diagnostic regions.
@@
For this Attn. variant, we apply standard attention [52] without any modifications, disregarding both scale and modality types.
@@
For hierarchical-scale message passing, we adopt the standard GraphSAGE [21] aggregation without any additional attention mechanism or modifications.
@@
To evaluate generalizability beyond TCGA datasets (BRCA, NSCLC, RCC), we further test HiVE-MIL and baselines on the Camelyon16 dataset [5], using the same experimental settings as described in the main paper. HiVE-MIL achieves the highest performance, outperforming the second-best model, ViLa-MIL, by 1.25% in Macro F1.
@@
ABMIL [27] 88.33 $\pm$4.44 92.97 $\pm$4.64 87.50 $\pm$4.67
DSMIL [33] 90.33 $\pm$2.77 94.95 $\pm$1.82 89.92 $\pm$2.85
CLAM-SB [38] 85.17 $\pm$14.16 85.07 $\pm$20.02 80.46 $\pm$22.48
DTFD-MIL (AFS) [38] 91.17 $\pm$4.25 93.35 $\pm$3.36 91.92 $\pm$4.43
ViLa-MIL [48] 92.33 $\pm$5.04 96.37 $\pm$2.28 91.97 $\pm$5.22
MSCPT [22] 88.50 $\pm$8.79 88.43 $\pm$13.44 88.29 $\pm$13.03
FOCUS [19] 90.83 $\pm$4.44 91.62 $\pm$6.52 90.23 $\pm$4.84
@@
We conduct an additional experiment to compare against MADELEINE [29], a representative WSI foundation model (not MIL-based), as it is pre-trained on CONCH patch features to generate slide-level embeddings. This setup ensures a fair comparison with our method (CONCH + HiVE-MIL).
@@
MADELEINE [29] (Linear Probing) 83.00 $\pm$4.10 90.30 $\pm$3.50 83.00 $\pm$4.10 80.40 $\pm$6.70 88.40 $\pm$6.40 80.30 $\pm$6.70
@@
To evaluate the robustness of HiVE-MIL to variations in textual input, we generate descriptions using four different LLMs: DeepSeek R1 [18], Grok 3 [53], Gemini 2.5 Pro [3], and GPT-4o [1]. The results in Table M.2 demonstrate that HiVE-MIL performs robustly across all LLMs. Although GPT-4o yields the best overall results, the performance differences among the LLMs are relatively small, indicating that HiVE-MIL does not heavily depend on a specific LLM. Importantly, regardless of the LLM used to generate textual descriptions, HiVE-MIL consistently outperforms all baseline methods. We expect performance to improve further as pathology-specific LLMs become available in the future, since the generated texts will be more accurate and less biased.
@@
LLM Model TCGA NSCLC TCGA BRCA
@@
DeepSeek R1 [18] 89.36 $\pm$3.88 95.77 $\pm$1.68 89.33 $\pm$3.92 85.60 $\pm$4.22 93.08 $\pm$2.43 85.49 $\pm$4.36
Grok 3 [53] 90.19 $\pm$2.33 96.33 $\pm$1.01 90.18 $\pm$2.33 85.80 $\pm$3.27 93.62 $\pm$1.80 85.71 $\pm$3.32
Gemini 2.5 Pro [3] 90.00 $\pm$2.52 96.47 $\pm$0.99 89.98 $\pm$2.55 86.65 $\pm$2.10 93.08 $\pm$1.79 86.61 $\pm$2.10
@@
WiKG [34] 891.82 0.0775 4630.56 1810.44 0.1539 8617.82
ViLa-MIL [48] 87.65 0.0242 528.20 116.00 0.0199 679.41
MSCPT [22] 860.21 0.1189 2074.89 894.17 0.1368 2695.92
FOCUS [19] 39.52 1.4500 353.21 46.20 1.9484 429.51

### v5

@@ 4.1 Experimental Settings
ABMIL [ABMIL] 84.30 ±2.22 90.97 ±0.60 84.28 ±2.21 81.04 ±3.05 87.50 ±5.38 80.93 ±3.04 88.43 ±1.95 96.17 ±0.76 86.95 ±2.33
DSMIL [DSMIL] 85.83 ±2.78 94.23 ±1.20 85.76 ±2.84 82.08 ±3.92 89.91 ±5.46 81.99 ±3.89 91.95 ±1.95 98.20 ±0.23 90.87 ±2.00
CLAM-SB [CLAM] 85.83 ±4.25 93.19 ±2.39 85.80 ±4.29 82.29 ±7.42 90.70 ±6.73 82.24 ±7.41 92.11 ±0.52 98.17 ±0.33 90.76 ±0.85
CLAM-MB [CLAM] 86.92 ±3.39 94.01 ±2.16 86.91 ±3.40 81.88 ±4.82 90.41 ±5.14 81.84 ±4.81 91.42 ±1.13 98.15 ±0.22 89.96 ±1.11
TransMIL [TransMIL] 85.90 ±3.36 93.38 ±2.11 85.88 ±3.36 82.50 ±5.37 89.69 ±4.54 82.38 ±5.36 89.27 ±2.34 97.75 ±0.69 87.66 ±2.95
DTFD-MIL [DTFD-MIL] 88.40 ±3.54 95.36 ±1.52 88.37 ±3.56 83.54 ±3.86 91.22 ±3.39 83.48 ±3.83 91.65 ±1.44 97.99 ±0.09 90.38 ±1.52
WiKG [WiKG] 82.24 ±3.13 91.17 ±1.62 82.15 ±3.21 79.58 ±6.17 87.42 ±6.54 79.44 ±6.39 89.73 ±2.37 97.65 ±0.67 87.84 ±3.12
ViLa-MIL [ViLa-MIL] 83.08 ±3.63 91.10 ±2.43 83.04 ±3.64 77.08 ±6.69 87.03 ±8.01 76.98 ±6.73 89.27 ±2.32 97.48 ±0.79 87.91 ±2.88
MSCPT [MSCPT] 80.06 ±5.20 88.06 ±6.28 79.95 ±5.24 79.79 ±8.22 87.33 ±6.78 79.69 ±8.21 92.03 ±1.52 98.03 ±0.35 90.89 ±1.94
FOCUS [MSCPT] 85.32 ±2.54 93.43 ±1.45 85.24 ±2.60 82.50 ±5.57 90.10 ±4.50 82.20 ±5.77 91.57 ±1.14 98.13 ±0.54 90.21 ±1.37
@@ 4.2 Main Results
We adopt the 16-shot evaluation setting as our main experimental setup, following [ViLa-MIL, MSCPT, MGPATH], and report results across three TCGA datasets (NSCLC, BRCA, RCC) and three vision-language pathology foundation models (PLIP, QuiltNet, CONCH). As shown in Table 1, HiVE-MIL consistently achieves the best performance across all settings, outperforming both traditional and VLM-based MIL methods, including their single-scale and multi-scale variants. Compared to the state-of-the-art baselines, HiVE-MIL achieves significant improvements across all datasets, with gains of up to +4.37% in ACC, +4.22% in AUC, and +4.14% in macro F1 on BRCA, and notable margins on NSCLC (up to +3.27% ACC, +3.28% AUC, +3.26% F1) and RCC (up to +1.07% ACC, +0.85% AUC, +1.64% F1). Even when paired with CONCH, the best pathology VLM to date, HiVE-MIL maintains consistent improvements across all datasets. Traditional MIL methods are based solely on visual features and large WSI labels, making them ineffective in the FSWC setting. Existing VLM-MIL methods, including multi-scale approaches, fail to capture hierarchical interactions across scales and do not effectively align modalities within the same scale. In contrast, HiVE-MIL explicitly models cross-scale hierarchies and intra-scale modality alignments, leading to superior performance.
@@ 5 Discussion
Acknowledgements.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) under grant numbers RS-2022-NR068758 and NRF-2025-25432820. We are also deeply grateful for the generous support provided by the Seegene Medical Foundation in South Korea.
@@ A.1 Dataset Descriptions
Following [ViLa-MIL], we split each dataset into training, validation, and test sets using a 4:3:3 ratio. For a few-shot evaluation, we repeat the random-splitting process five times and report averaged results. Due to the class imbalance in the TCGA BRCA (approximately 20:80), we balance the test set during sampling to avoid misleading final metric results. All details regarding the data splits used are provided in the splits directory of our GitHub repository (linked in the Abstract).
@@ A.2 Hierarchical Patch Extraction
Existing MIL methods [DSMIL, ViLa-MIL, MSCPT, MGPATH] typically process patch features at different scales independently, without preserving any hierarchical structure between them. This design hinders the ability to model structured cross-scale interactions, which are essential in tasks such as WSI classification, where coarse tissue-level patterns and fine-grained cellular features are often interdependent. To address this, we propose a hierarchical extraction pipeline with an explicit 1-to-16 mapping across scales: each low-scale ($5\times$) patch corresponds to a spatial region at the high scale ($20\times$), which is uniformly subdivided into a fixed $4\times 4$ grid, yielding 16 high-scale patches. This structured mapping ensures that each high-scale patch can be directly and uniquely traced back to its low-scale parent, enabling consistent alignment and hierarchical reasoning across scales.
@@ B.1 Class-Discriminative Relevance
We evaluate the faithfulness of the generated descriptions using the LLM-as-a-Judge [LLM_as_Judge], where LLMs are prompted to infer class labels based solely on the generated descriptions. We employ two evaluators: GPT-4o and GPT o3, to assess descriptions generated by several LLMs: Deep Seek R1 [deepseek-r1], Grok 3 [grok3], Gemini 2.5 Pro [gemini], and GPT-4o [GPT-4o]. Although not perfect, the number of correct predictions is still consistently high, suggesting that the generated text is generally reliable while capturing class-discriminative signals.
@@
DeepSeek R1 [deepseek-r1] 32 (8/24) 28 (8/20) 28 (8/20) 31 (8/23) 28 (8/20)
Grok 3 [grok3] 32 (8/24) 28 (7/21) 28 (8/20) 32 (8/24) 21 (8/23)
Gemini 2.5 Pro [grok3] 32 (8/24) 30 (8/22) 29 (7/22) 30 (7/23) 32 (8/24)
GPT-4o [GPT-4o] 32 (8/24) 30 (8/22) 30 (7/23) 31 (8/23) 32 (8/24)
@@ B.2 Morphology- and Concept-Level Semantic Validation
Table B.2 reports intra- and inter-category text similarity scores, where intra denotes similarities among descriptions of the same morphology or concept, and inter denotes similarities across different ones. Similarity is computed separately at each scale and jointly across both scales using the CONCH text encoder [CONCH] and cosine similarity. These results demonstrate that the descriptions encode highly discriminative and consistent semantics across both morphological groups and scales, and across different LLMs.
@@
DeepSeek R1 [deepseek-r1] 1.000 (0.157) 0.515 (0.164) 0.456 (0.158) 1.000 (0.165) 0.438 (0.141) 0.412 (0.138)
Grok 3 [grok3] 1.000 (0.123) 0.511 (0.170) 0.434 (0.160) 1.000 (0.186) 0.490 (0.214) 0.415 (0.212)
Gemini 2.5 Pro [gemini] 1.000 (0.111) 0.558 (0.165) 0.478 (0.150) 1.000 (0.206) 0.491 (0.127) 0.426 (0.142)
GPT-4o [GPT-4o] 1.000 (0.066) 0.516 (0.114) 0.454 (0.109) 1.000 (0.163) 0.477 (0.108) 0.413 (0.125)
@@ Appendix E Message Passing Details
The message passing process is as follows. During neighbor aggregation, information is collected from neighboring nodes connected via a specific relation type (i.e., edge type). These include intra-scale relations, such as valid patch-text connections at low or high scale, and hierarchical relations, such as connections between low-scale and high-scale patches, or between low-scale and high-scale text. The aggregated neighbor features are then concatenated with the node’s own features and passed through a relation-specific linear transformation (i.e., weight matrix). For intra-scale relations, we adopt the GraphSAGE [GraphSAGE] operator via the HeteroConv module, whereas for hierarchical relations, we employ a Modality-Scale Attention (MSA) mechanism that accounts for both scale directionality and modality type. Finally, a non-linear activation function (ReLU) is applied to enhance representational capacity. By decoupling both aggregation and transformation parameters by relation type, our model effectively encodes heterogeneous and hierarchical information.
@@ Appendix H Implementation Details
We set the logit scaling factor $\gamma$ to 4.5871 (PLIP [PLIP]), 4.6052 (QuiltNet [QuiltNet]), and 4.0315 (CONCH [CONCH]), using the values provided by each pre-trained VLM. We retain the top-2 patches at $5\times$ and the top-100 patches at $20\times$ scales to compute the final logits. Pooling- and MIL-based methods use only the image encoder, while VLM-based MIL methods leverage both image and text encoders. Most methods operate on single-scale $20\times$ patches, whereas DSMIL [DSMIL], ViLa-MIL [ViLa-MIL], MSCPT [MSCPT], and HiVE-MIL (Ours) utilize multi-scale inputs from both $5\times$ and $20\times$ patches.
Our implementation is based on the official ViLa-MIL codebase [ViLa-MIL] and all baselines as well as HiVE-MIL are implemented within this unified framework to ensure fair comparison. Our graph-based modules are implemented using PyTorch Geometric [PyTorch_Geometric]. All experiments are conducted on Ubuntu 20.04.6 using a workstation equipped with two NVIDIA A100 GPUs (40 GB each); however, only one GPU is used for training each model. Complete package versions and dependencies are listed in the requirements.txt file available in our GitHub repository (linked in the Abstract).
@@ Appendix I Baselines
  • PLIP [PLIP]: A vision-language pathology model pretrained on OpenPath, a large dataset of 208,414 image–text pairs curated from public platforms like medical Twitter, enabling strong zero-shot classification and case retrieval via image or language queries.
  • QuiltNet [QuiltNet]: A vision-language model pretrained on Quilt, a 1M-pair dataset curated from YouTube, medical Twitter, and academic sources, enabling strong zero-shot classification and cross-modal retrieval across diverse histopathology datasets.
  • CONCH [CONCH]: A recent state-of-the-art vision-language pathology foundation model, pretrained on 1.17 million image–caption pairs from diverse biomedical sources, enabling strong zero-shot and transferable performance across 14 benchmarks spanning classification, segmentation, captioning, and retrieval tasks.
@@
  • ABMIL [ABMIL]: An attention-based pooling framework that assigns adaptive weights to instances for more informative bag-level aggregation.
  • DSMIL [DSMIL]: A dual-stream MIL model where one stream identifies a critical instance via max pooling, while the other aggregates instances by distance-weighted similarity.
  • CLAM [CLAM]: A weakly-supervised MIL framework that combines attention pooling with instance-level clustering to enhance interpretability and slide-level prediction.
  • TransMIL [TransMIL]: A transformer-based MIL model that captures inter-instance correlations by modeling both spatial and morphological relationships among patches.
  • DTFD-MIL [DTFD-MIL]: A double-tier MIL framework that introduces pseudo-bags to improve training diversity and derives instance probabilities under an attention-based setting.
  • WiKG [WiKG]: A graph-based MIL approach that constructs WSIs as knowledge graphs with directed edges and updates features using knowledge-aware attention.
  • ViLa-MIL [ViLa-MIL]: A dual-scale VLM-MIL model generates LLM-based prompts and uses a prototype-guided image decoder and a context-guided text decoder.
  • MSCPT [MSCPT]: A multi-scale VLM MIL model that integrates multi-scale visual inputs with LLM-generated prompts through graph-based reasoning and cross-scale aggregation.
  • FOCUS [FOCUS]: A single-scale VLM-MIL model that integrates pathology foundation models with language priors for focused analysis of diagnostic regions.
@@
For this Attn. variant, we apply standard attention [MHA] without any modifications, disregarding both scale and modality types.
@@
For hierarchical-scale message passing, we adopt the standard GraphSAGE [GraphSAGE] aggregation without any additional attention mechanism or modifications.
@@
To evaluate generalizability beyond TCGA datasets (BRCA, NSCLC, RCC), we further test HiVE-MIL and baselines on the Camelyon16 dataset [Camelyon16], using the same experimental settings as described in the main paper. HiVE-MIL achieves the highest performance, outperforming the second-best model, ViLa-MIL, by 1.25% in Macro F1.
@@
ABMIL [ABMIL] 88.33 $\pm$4.44 92.97 $\pm$4.64 87.50 $\pm$4.67
DSMIL [DSMIL] 90.33 $\pm$2.77 94.95 $\pm$1.82 89.92 $\pm$2.85
CLAM-SB [CLAM] 85.17 $\pm$14.16 85.07 $\pm$20.02 80.46 $\pm$22.48
DTFD-MIL (AFS) [CLAM] 91.17 $\pm$4.25 93.35 $\pm$3.36 91.92 $\pm$4.43
ViLa-MIL [ViLa-MIL] 92.33 $\pm$5.04 96.37 $\pm$2.28 91.97 $\pm$5.22
MSCPT [MSCPT] 88.50 $\pm$8.79 88.43 $\pm$13.44 88.29 $\pm$13.03
FOCUS [FOCUS] 90.83 $\pm$4.44 91.62 $\pm$6.52 90.23 $\pm$4.84
@@
We conduct an additional experiment to compare against MADELEINE [Madeleine], a representative WSI foundation model (not MIL-based), as it is pre-trained on CONCH patch features to generate slide-level embeddings. This setup ensures a fair comparison with our method (CONCH + HiVE-MIL).
@@
MADELEINE [Madeleine] (Linear Probing) 83.00 $\pm$4.10 90.30 $\pm$3.50 83.00 $\pm$4.10 80.40 $\pm$6.70 88.40 $\pm$6.40 80.30 $\pm$6.70
@@
To evaluate the robustness of HiVE-MIL to variations in textual input, we generate descriptions using four different LLMs: DeepSeek R1 [deepseek-r1], Grok 3 [grok3], Gemini 2.5 Pro [gemini], and GPT-4o [GPT-4o]. The results in Table M.2 demonstrate that HiVE-MIL performs robustly across all LLMs. Although GPT-4o yields the best overall results, the performance differences among the LLMs are relatively small, indicating that HiVE-MIL does not heavily depend on a specific LLM. Importantly, regardless of the LLM used to generate textual descriptions, HiVE-MIL consistently outperforms all baseline methods. We expect performance to improve further as pathology-specific LLMs become available in the future, since the generated texts will be more accurate and less biased.
@@
LLM TCGA NSCLC TCGA BRCA
@@
DeepSeek R1 [deepseek-r1] 89.36 $\pm$3.88 95.77 $\pm$1.68 89.33 $\pm$3.92 85.60 $\pm$4.22 93.08 $\pm$2.43 85.49 $\pm$4.36
Grok 3 [grok3] 90.19 $\pm$2.33 96.33 $\pm$1.01 90.18 $\pm$2.33 85.80 $\pm$3.27 93.62 $\pm$1.80 85.71 $\pm$3.32
Gemini 2.5 Pro [gemini] 90.00 $\pm$2.52 96.47 $\pm$0.99 89.98 $\pm$2.55 86.65 $\pm$2.10 93.08 $\pm$1.79 86.61 $\pm$2.10
@@
WiKG [WiKG] 891.82 0.0775 4630.56 1810.44 0.1539 8617.82
ViLa-MIL [ViLa-MIL] 87.65 0.0242 528.20 116.00 0.0199 679.41
MSCPT [MSCPT] 860.21 0.1189 2074.89 894.17 0.1368 2695.92
FOCUS [FOCUS] 39.52 1.4500 353.21 46.20 1.9484 429.51